OpenStudy (anonymous):

The point (2,3) is on the graph y=f(x). Determine the corresponding coordinates of this point on the graph y=-2(f(2(x+5)))-4

OpenStudy (anonymous):

sure can use some help

OpenStudy (loser66):

To me, I would change a little bit from the notation of the second function to make it easily to follow . Let say A = -2 f(B) - 4 the first function is the parent function which give us 3 = f(2) now find the relationship between them by finding A respect to f(B) , in this case, find A of f(2) A = -2f(2) -4 = -2 *3 -4 = -10, So, the new y (which is A) is -10 f(B) = f(2) = f(2(x + 5) so, 2(x+5) =2 --> x = -4 therefore, in the new function, the image of (2,3) is (-4,-10)

OpenStudy (anonymous):

w8 did u just use substitution?